OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

y=-3x + 12

OpenStudy (anonymous):

mad smart could u help me with this..... Choose the equation of the line passing through the point (1, -5) and perpendicular to y = 1/5x - 10. y = 5x + 10 y = 5x y = -5x - 10 y = -5x

OpenStudy (ajprincess):

If m is the slope of a line. then the slope of the line perpendicular to the line will be -1/m.  \[m=\frac{y-y_1}{x-x_1}\]  Is that clear?